# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

### v2

@@ 3.2.1 Stage 1: Causal Visual Encoding
We instantiate $E_{cause}$ and $E_{effect}$ as two separate instances of the CLIP-ViT (ViT-B/32) model Radford et al. (2021), adapted from CLIP4Clip Luo et al. (2022) (see Appendix A.6 for more details). For both Cause and Effect text encoders ($T_{cause}$ and $T_{effect}$ respectively), we employ a 12-layer transformer with a hidden size of $512$ and $8$ attention heads, using the weights from the pre-trained CLIP Radford et al. (2021) text encoder. As in CLIP Radford et al. (2021), the [EOS] token’s representation from the last layer of the text encoders is used as the feature representation for the input text. Similar to CLIP4Clip, we adopt cosine similarity to measure the similarity between the video embedding (mean pooled) and the text embedding. Given a video embedding $\hat{r}_{i}$ and a text embedding $t_{j}$, the similarity function is defined as $s(\hat{r}_{i},t_{j})=\frac{t_{j}^{\top}\hat{r}_{i}}{|t_{j}||\hat{r}_{i}|}$. During training, each model is optimized using a contrastive loss, as in CLIP Radford et al. (2021), over its respective cause/effect portion from CTN captions. For a batch of $N$ video-cause text pairs $(v_{i},c_{i})$ and video-effect text pairs $(v_{i},e_{i})$, the contrastive losses for the cause and effect video encoders are defined as:
@@ 3.2.2 Stage 2: CTN Caption Generation
where $\hat{y}_{<t}$ denotes the previously generated words. For training and evaluation, we combine the cause and effect parts of the ground-truth CTN caption with a space in between to form a single, coherent caption. This combined caption serves as the target for the decoder during training and the reference for evaluating the generated captions using diffferent evaluation metrics. The entire network is trained end-to-end using the cross-entropy loss between the generated caption $\hat{y}$ and the ground-truth CTN caption $y$:
@@ 4 Experiments
We conduct extensive experiments to validate our proposed CEN architecture and the new CTN video captioning benchmark. This section details the datasets used, the experimental setup, and the evaluation metrics. We also present comprehensive results and evaluation against the state-of-the-art video captioning methods and vision-language models (VLMs) demonstrating the performance of our network on the CTN benchmark.
4.1 Datasets
We evaluate our CTN captions and CEN on two widely-used video captioning datasets: 1) MSR-VTT Xu et al. (2016), consists of 10,000 videos with 200,000 human-annotated captions, and MSVD Chen & Dolan (2011), with 1,970 videos focused on human activities with 78,800 captions. The captions in these datasets lack explicit representations of the causal relationships, and temporal dynamics which we address by generating CTN captions as detailed in Section 3.1.
@@ 4.2 Evaluation Metrics
For quantifying the performance of our method CEN against the SOTA methods, we use three metrics as per Celikyilmaz et al. (2020); CIDEr (C) Vedantam et al. (2015): Measures alignment of generated captions with references by highlighting frequently occurring terms, capturing the ability to reproduce salient causal-temporal narrative elements. ROUGE-L (R-L) Lin (2004): Evaluates the longest common subsequence between generated and reference captions, considering both precision and recall for assessing semantic similarity, including temporal dynamics and causal relationships. SPICE (S) Anderson et al. (2016): Evaluates semantic quality by considering the overlap of scene graphs between generated and reference captions, effectively assessing the ability to capture and express underlying causal relationships and event sequences.
@@ 4.3 Implementation Details
We generate the CTN captions using the Mixtral of Experts LLM Jiang et al. (2024), running on A100-80GB GPUs. All models, including our proposed CEN architecture, the SOTA models (SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023), and GIT Wang et al. (2022)) and the VLMs VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) are implemented using PyTorch. For the SOTA models, we follow the hyperparameter settings specified in their respective methods for training on the MSR-VTT and MSVD datasets. Our CEN model is trained using the Adam optimizer with learning rates of $1\times 10^{-4}$ (stage 1) and $1\times 10^{-6}$ (stage 2) and a batch size of 64 for 10 epochs (stage 1) and 50 epochs (stage 2). For comparison with recent Vision-Language Models (VLMs), we fine-tune VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) using both LoRA and simple fine-tuning approaches on our CTN benchmark. We use the recommended hyperparameters for each model during fine-tuning. Further details are provided in the Appendix A.1.
@@ 4.4 Results and Discussion
We comprehensively evaluate the performance of our proposed CEN architecture against several SOTA methods, including SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023), and GIT Wang et al. (2022) and the VLMs VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) on our generated CTN captions benchmark datasets.
Table 1: Comparison of our CEN architecture against SOTA methods and VLMs on the MSVD and MSR-VTT datasets. The best results in each category are in bold. R-L, C, and S denote ROUGE-L, CIDEr, and SPICE scores, respectively. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) SEM-POS 25.39 37.16 14.46 20.11 26.01 12.09 AKGNN 25.11 35.08 14.55 21.42 25.90 11.99 GIT 27.51 45.63 15.58 24.51 32.43 13.70 VideoLLaVA (Zero-shot) 21.80 30.55 14.67 19.33 16.24 12.49 VideoLLaVA (LoRA FT) 24.56 34.98 15.41 21.21 18.97 13.28 VideoLLaVA (Simple FT) 25.61 36.12 16.09 22.18 19.98 13.07 ShareGPT4Video (Zero-shot) 21.66 27.06 14.06 20.27 17.08 12.21 ShareGPT4Video (LoRA FT) 24.39 30.72 14.83 22.09 19.83 13.02 ShareGPT4Video (Simple FT) 25.32 31.67 14.92 23.01 20.76 13.28 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76 Table 2: Ablation study results on the MSVD and MSR-VTT datasets. $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$ are variants of our CEN architecture, while Zero Shot v.v. and Fine-tune v.v. represent cross-dataset evaluation settings. The best results in each category are in bold. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) $\text{E}_{\text{JointCE}}$ 30.93 55.72 17.04 27.34 45.97 15.07 Only $\text{E}_{\text{cause}}$ 30.72 56.42 18.43 27.19 47.10 15.21 Only $\text{E}_{\text{effect}}$ 30.70 57.14 17.89 27.24 45.58 15.19 Zero Shot v.v. 27.16 39.65 14.45 24.73 29.76 12.26 Fine-tune v.v. 31.78 65.60 19.39 27.47 47.74 15.65 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
Figure 4: (a) UMAP visualization of video features learned from CTN (red) and original (blue) captions on MSR-VTT, showing non-overlapping feature spaces. (b) UMAP visualization of video features learned from cause (black) and effect (orange) parts of CTN captions on MSR-VTT, showing near-complete overlap.
@@ 4.4.1 Quantitative Results
CEN outperformed all SOTA methods and the VLMs across all metrics and datasets as shown in Table 2. On MSVD, CEN surpassed the next best (GIT) by 3.95 ROUGE-L, 17.88 CIDEr, and 3.67 SPICE points. On MSR-VTT, CEN led GIT by 3.39 ROUGE-L, 17.44 CIDEr, and 2.06 SPICE points. These significant gains highlight CEN’s effectiveness in capturing causal narratives and temporal dynamics.
We compare CEN’s performance against zero-shot and fine-tuned (FT) versions of VideoLLaVA and ShareGPT4Video. CEN consistently outperforms these models, even after fine-tuning, demonstrating the effectiveness of our specialized architecture for causal-temporal narrative generation. For example, on MSR-VTT, CEN achieves a CIDEr score of 49.87, compared to 19.98 for VideoLLaVA (Simple FT) and 20.76 for ShareGPT4Video (Simple FT).
To demonstrate generalization capability, we perform cross-dataset evaluations (Table 2). CEN shows strong zero-shot performance (Zero Shot v.v.) and significant improvements after fine-tuning (Fine-tune v.v.), outperforming models trained directly on the target datasets. For instance, when trained on MSR-VTT and tested on MSVD, CEN achieves a CIDEr score of 65.60 after fine-tuning, compared to 45.63 for GIT trained directly on MSVD.
@@ 4.4.2 Qualitative Results
Qualitative examples, in Figure 5, demonstrate CEN’s strength in accurately articulating causal relationships and event sequences, while SOTA methods and recent VLMs struggle. For example: (a) Video game scene: CEN accurately captures the causality ("performs a fatality move" causing "killing another"), while other methods miss this crucial relationship. SEM-POS and GIT provide overly simplistic descriptions, while AKGNN generates irrelevant details. VLMs (Video-LLaVA and ShareGPT4Video) misinterpret the scene entirely.
(b) Paper folding: CEN correctly identifies both the cause (folding paper) and effect (creating a paper airplane). Other methods either miss the causal relationship or misinterpret the action (e.g., ShareGPT4Video’s "small square" instead of an airplane).
(c) Soccer scene: CEN accurately links the precise kick to the goal, capturing both cause and effect. Other methods either focus on a single action or provide generic descriptions of a soccer game.
(d) Singing performance: CEN captures the boy’s decision to perform (cause) and the audience’s reaction (effect). Other methods mostly describe the act of singing without the causal-temporal context.
@@
4.4.3 Ablation Results
To examine the effectiveness of CEN, we compare it with 4 baselines in Table 2:
a) $\text{E}_{\text{JointCE}}$: Instead of using two separate video encoders for cause and effect, we train one CLIP-ViT using the combined cause and effect captions separated by space for the text encoder. The performance drops across all metrics and datasets compared to the CEN architecture, which underscores the benefits of dedicated encoders for capturing cause and effect dynamics.
b) Only Ecause and Only Eeffect: These ablations employ only the cause encoder or only the effect encoder, respectively. While retaining reasonable performance, both variants fall short of the performance of the full model across all metrics and datasets.
c) Zero Shot v.v.: In this zero-shot setting, we evaluate the models trained on one dataset (MSVD or MSR-VTT) against the other dataset, without any fine-tuning on the other dataset. The substantial performance drop highlights the importance of dataset-specific fine-tuning for effective causal-temporal narrative learning.
d) Fine-tune v.v.: Building upon the zero-shot case, this variant involves fine-tuning the models trained on one dataset (MSVD or MSR-VTT) against the other dataset. Notably, this fine-tuning process leads to improved performance on the MSVD dataset, even surpassing the CEN model. This observation demonstrates the potential for transfer learning from larger datasets and the ability to leverage their knowledge effectively on smaller datasets through fine-tuning.
Overall, the ablation study validates the efficacy of our proposed CEN architecture, emphasizing the significance of dedicated encoders for capturing cause and effect relationships independently. The performance drops observed in $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$, and Only $\text{E}_{\text{effect}}$ results, compared to the CEN model, clearly indicate the necessity of explicitly embedding causal-temporal narrative understanding within the architecture through dedicated cause and effect encoders.

### v3

@@ 3.2.1 Stage 1: Causal Visual Encoding
We instantiate $E_{cause}$ and $E_{effect}$ as two separate instances of the CLIP-ViT (ViT-B/32) model Radford et al. (2021), adapted from CLIP4Clip Luo et al. (2022) (see Appendix A.6 for more details). For both Cause and Effect text encoders ($T_{cause}$ and $T_{effect}$ respectively), we employ a 12-layer transformer with a hidden size of $512$ and $8$ attention heads, using the weights from the pre-trained CLIP Radford et al. (2021) text encoder. As in CLIP Radford et al. (2021), the [EOS] token’s representation from the last layer of the text encoders is used as the feature representation for the input text. Similar to CLIP4Clip, we use mean pooling for the video embedding and then, adopt cosine similarity to measure the similarity between the video embedding (mean pooled) and the text embedding. Given a video embedding $\hat{r}_{i}$ and a text embedding $t_{j}$, the similarity function is defined as $s(\hat{r}_{i},t_{j})=\frac{t_{j}^{\top}\hat{r}_{i}}{|t_{j}||\hat{r}_{i}|}$. During training, each model is optimized using a contrastive loss, as in CLIP Radford et al. (2021), over its respective cause/effect portion from CTN captions. For a batch of $N$ video-cause text pairs $(v_{i},c_{i})$ and video-effect text pairs $(v_{i},e_{i})$, the contrastive losses for the cause $(\mathcal{L}_{cause})$ and effect $(\mathcal{L}_{effect})$ video encoders are defined in 1 and 2 respectively:
@@ 3.2.2 Stage 2: CTN Caption Generation
where $\hat{y}_{<t}$ denotes the previously generated words. For training and evaluation, we combine the cause and effect parts of the ground-truth CTN caption with a space in between to form a single, coherent caption. We opt for space separation to maintain consistency with the raw input format of the training datasets (MSRVTT and MSVD), which contain no separators and also, for a fair comparison with the SOTA. This combined caption serves as the target for the decoder during training and the reference for evaluating the generated captions using diffferent evaluation metrics. The entire network is trained end-to-end using the cross-entropy loss between the generated caption $\hat{y}$ and the ground-truth CTN caption $y$:
@@ 4 Experiments
We conduct extensive experiments to validate our proposed CEN architecture on the new CTN video captioning benchmark dataset. We present comprehensive evaluation against the state-of-the-art video captioning methods and vision-language models (VLMs) demonstrating the superior performance of the proposed network on the CTN benchmark dataset.
@@ 4.1 Evaluation Metrics
For quantitative comparison, we use three metrics as per Celikyilmaz et al. (2020); CIDEr (C) Vedantam et al. (2015); Shi et al. (2022): Measures alignment of generated captions with references by highlighting frequently occurring terms, capturing the ability to reproduce salient causal-temporal narrative elements. ROUGE-L (R-L) Lin (2004): Evaluates the longest common subsequence between generated and reference captions, considering both precision and recall for assessing semantic similarity, including temporal dynamics and causal relationships. SPICE (S) Anderson et al. (2016): Evaluates semantic quality by considering the overlap of scene graphs between generated and reference captions, effectively assessing the ability to capture causal relationships and event sequences.
@@ 4.2 Implementation Details
We generate the CTN captions using the Mixtral of Experts LLM Jiang et al. (2024), running on A100-80GB GPUs. For the SOTA models, we follow the hyperparameter settings specified in their respective methods for training on the MSR-VTT and MSVD datasets. Our CEN model is trained using the Adam optimizer with learning rates of $1\times 10^{-4}$ (stage 1) and $1\times 10^{-6}$ (stage 2) and a batch size of 64 for 10 epochs (stage 1) and 50 epochs (stage 2). For comparison with recent Vision-Language Models (VLMs), we fine-tune VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) using both LoRA and simple fine-tuning approaches on our CTN benchmark dataset. We use the recommended hyperparameters for each model during fine-tuning. Further details are provided in the Appendix A.1.
@@ 4.3 Results and Discussion
We comprehensively evaluate the performance of our proposed CTN generation pipeline and also, the CEN architecture against several SOTA methods, including SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023), and GIT Wang et al. (2022) and the VLMs VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) on our generated CTN captions benchmark datasets.
@@ 4.3.1 Quantitative Results
CTN Captions Benchmark Dataset: We generate our CTN captions (1 caption per video) using two widely-used video captioning datasets: 1) MSRVTT Xu et al. (2016), consists of 10,000 videos with 20 human-annotated captions per video, and MSVD Chen & Dolan (2011), with 1,970 videos focused on human activities with approx. 50 captions per video. The length of the caption is on avg. 19 words (cause=10 words, effect=9 words) for MSRVTT-CTN and 17 words (cause=8 words, effect=7 words) for MSVD-CTN. Then, we train as well as test our CEN, SOTA methods and VLMs on the MSVRTT-CTN and MSVD-CTN.
Our CTN generation approach generates good quality captions with 52.1% CTN captions exceeding 0.27 EMScore (threshold=0.2). Then, in the human evaluation study, the overall mean quality score of 4.8 ($\sigma$ = 0.40) on a 5-point Likert scale, indicates high performance across all criteria; 93% of captions receive scores of 4 or higher across all three dimensions, with 84% achieving perfect scores. To ensure reliability across raters, we calculate the intraclass correlation coefficient (ICC). The ICC for absolute agreement among raters is 0.87 (95% CI: 0.83-0.91), indicating high inter-rater reliability. These results validate our automatic generation and evaluation process, demonstrating that our CTN captions benchmark dataset provides high-quality, coherent representations of causal-temporal narratives in videos. Further details are provided in Appendix A.5.
Table 1: Comparison of our CEN architecture against SOTA methods and VLMs on the MSVD-CTN and MSRVTT-CTN datasets. The best results in each category are in bold. R-L, C, and S denote ROUGE-L, CIDEr, and SPICE scores, respectively.
Method MSVD-CTN MSRVTT-CTN
R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ )
SEM-POS 25.39 37.16 14.46 20.11 26.01 12.09
AKGNN 25.11 35.08 14.55 21.42 25.90 11.99
GIT 27.51 45.63 15.58 24.51 32.43 13.70
VideoLLaVA (Zero-shot) 21.80 30.55 14.67 19.33 16.24 12.49
VideoLLaVA (LoRA FT) 24.56 34.98 15.41 21.21 18.97 13.28
VideoLLaVA (Simple FT) 25.61 36.12 16.09 22.18 19.98 13.07
ShareGPT4Video (Zero-shot) 21.66 27.06 14.06 20.27 17.08 12.21
ShareGPT4Video (LoRA FT) 24.39 30.72 14.83 22.09 19.83 13.02
ShareGPT4Video (Simple FT) 25.32 31.67 14.92 23.01 20.76 13.28
CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
Figure 4: (a) UMAP visualization of video features learned from CTN (red) and original (blue) captions on MSR-VTT, showing non-overlapping feature spaces. (b) UMAP visualization of video features learned from cause (black) and effect (orange) parts of CTN captions on MSRVTT-CTN, showing near-complete overlap.
Cause-Effect Network (CEN): CEN outperformed all SOTA methods and the VLMs across all metrics and datasets as shown in Table 1. On MSVD-CTN, CEN surpassed the next best (GIT) by 3.95 ROUGE-L, 17.88 CIDEr, and 3.67 SPICE points. On MSRVTT-CTN, CEN led GIT by 3.39 ROUGE-L, 17.44 CIDEr, and 2.06 SPICE points. These significant gains highlight CEN’s effectiveness in capturing causal narratives and temporal dynamics.
We compare CEN’s performance against zero-shot and fine-tuned (FT) versions of VideoLLaVA and ShareGPT4Video. CEN consistently outperforms these models, even after fine-tuning, demonstrating the effectiveness of our specialized architecture for causal-temporal narrative generation. For example, on MSRVTT-CTN, CEN achieves a CIDEr score of 49.87, compared to 19.98 for VideoLLaVA (Simple FT) and 20.76 for ShareGPT4Video (Simple FT).
@@
Figure 5: Ablation study results on the MSVD-CTN and MSRVTT-CTN datasets. $\text{E}_{\text{combined}}$, w/o FT CLIPs, Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$ are baselines of our CEN architecture, while Zero Shot X and Fine-tune X represent cross-dataset evaluation settings. The best results in each category are in bold.
Method MSVD-CTN MSRVTT-CTN
R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ )
$\text{E}_{\text{combined}}$ 30.93 55.72 17.04 27.34 45.97 15.07
w/o FT - single CLIP 27.81 46.23 15.82 25.34 32.31 14.27
w/o FT - Two CLIP 28.40 53.84 16.58 26.10 40.92 14.55
Only $\text{E}_{\text{cause}}$ 30.72 56.42 18.43 27.19 47.10 15.21
Only $\text{E}_{\text{effect}}$ 30.70 57.14 17.89 27.24 45.58 15.19
Zero Shot X 27.16 39.65 14.45 24.73 29.76 12.26
Fine-tune X 31.78 65.60 19.39 27.47 47.74 15.65
CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
Figure 6: EMScore distribution of initial CTN captions showing 66.4% exceed quality threshold $\theta$=0.2
4.3.2 Ablation Results
CTN Captions Benchmark Dataset: Figure 6 demonstrates the distribution of caption quality without automatic evaluation filtering. The EMScore distribution across 11,970 videos shows 66.4% of initially generated captions achieve scores above $\theta$=0.2, with a peak around EMScore=0.24. The remaining 33.6% of captions fall below our quality threshold, highlighting the importance of our automatic evaluation and regeneration process in maintaining CTN quality.
Cause-Effect Network (CEN): To examine the effectiveness of CEN, we compare it with 6 baselines in Table 6:
a) $\text{E}_{\text{combined}}$: Instead of using two separate video encoders for cause and effect, we train one CLIP-ViT using the combined cause and effect captions separated by space for the text encoder. The performance drops across all metrics and datasets compared to the CEN architecture, which underscores the benefits of dedicated encoders for capturing cause and effect dynamics.
b) w/o FT - Single CLIP: This ablation is performed using one CLIP encoder and no Fine-tuning (FT) on cause, effect, or cause+effect (combined). The results show the effectiveness of CTN fine-tuning.
c) w/o FT - Two CLIP: This ablation is performed using two CLIP $\text{E}_{\text{cause}}$ and $\text{E}_{\text{effect}}$ encoders (as in our CEN) without performing the Stage 1 fine-tuning (FT) on cause and effect. This demonstrates the effectiveness of CTN fine-tuning and also, the performance increase in comparison to "w/o FT - Single CLIP" demonstrates the effectiveness of CEN.
d) Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$: These ablations employ only the cause encoder or only the effect encoder, respectively. While retaining reasonable performance, both variants fall short of the performance of the full model across all metrics and datasets.
e) Zero Shot X: In this zero-shot setting, we evaluate the models trained on one dataset (MSVD-CTN or MSRVTT-CTN) against the other dataset, without any fine-tuning on the other dataset. The results are either better than or comparable with SOTA and VLMs in Table 1.
f) Fine-tune X: Similar to the Zero Shot X, this variant involves fine-tuning the models trained on cross-datasets. Notably, this fine-tuning process leads to improved performance on the MSVD dataset. This observation demonstrates the potential for transfer learning of CEN model from larger datasets i.e. MSRVTT-CTN and the ability to leverage their knowledge effectively on smaller datasets i.e. MSVD-CTN through fine-tuning Lin et al. (2022); Ventura et al. (2024)
Overall, the ablation study validates the efficacy of our proposed CEN architecture and approach, emphasizing the significance of dedicated encoders for capturing cause and effect relationships independently as the performance drops are observed in $\text{E}_{\text{combined}}$, w/o FT - Single CLIP, w/o FT - Two CLIP, Only $\text{E}_{\text{cause}}$, and Only $\text{E}_{\text{effect}}$ baselines. We also provide LLM-based evaluations in Appendix A.9.
@@ 4.3.3 Qualitative Results
Qualitative examples, in Figure 7, demonstrate CEN’s strength in accurately articulating causal relationships and event sequences, while SOTA methods and recent VLMs struggle. For example: (a) Video game scene: CEN accurately captures the causality ("performs a fatality move" causing "killing another"), while other methods miss this crucial relationship. SEM-POS and GIT provide overly simplistic descriptions, while AKGNN generates irrelevant details. VLMs (Video-LLaVA and ShareGPT4Video) misinterpret the scene entirely.; (b) Paper folding: CEN correctly identifies both the cause (folding paper) and effect (creating a paper airplane). Other methods either miss the causal relationship or misinterpret the action (e.g., ShareGPT4Video’s "small square" instead of an airplane).; (c) Soccer scene: CEN accurately links the precise kick to the goal, capturing both cause and effect. Other methods either focus on a single action or provide generic descriptions of a soccer game.; (d) Singing performance: CEN captures the boy’s decision to perform (cause) and the audience’s reaction (effect). Other methods mostly describe the act of singing without the causal-temporal context.